# SGFormer: Spherical Geometry Transformer for 360 Depth Estimation

### v2

@@ IV-B Comparison Results
Quantitative Analysis: To validate the effectiveness of our approach, we compare our SGFormer model with the current state-of-the-art methods, including strategies for bi-projection fusion and long-range dependencies. As shown in Table I, our model ranks first in nearly all evaluation metrics, indicating superior overall performance. Specifically, on the one hand, compared with Bifuse [1] based on bi-projection fusion, our model has achieved a 53% improvement (Abs.rel) and 41% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is significantly improved by 66% (Abs.rel) and 51% (RMS.lin). On the other hand, compared with EGFormer [2] based on long-range dependencies, our model achieves an 11% improvement (Abs.rel) and 12% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, SGFormer achieves 12% (Abs.rel) and 9% (RMS.lin) gain. In quantitative comparison, we achieve significant performance improvements because we fully leverage the geometric priors of the sphere to relieve the panoramic distortion. The lower quality of depth estimation in SliceNet [12] is mainly due to the detail loss when processing equirectangular images in the vertical direction.
@@
More Comparison Results: We also compare our method with the latest methods from 2024 on additional datasets. For a fair comparison, we strictly follow the Elite360D [43]. The quantitative comparison is shown in Table II, we achieve better performance on the Stanford2D3D dataset [41] and Matterport3D dataset [40].
TABLE III: Ablation studies of individual components. We begin with a basline model, adding each component sequentially. We conduct a series of ablation studies on the S3D (Structured3D) and Pano3D datasets. Results in bold indicate the best results. (CLE: curve local embedding, GCPE: global conditional position embedding, CR: circular rotation, BRP: bipolar re-projection)
@@ IV-C Ablation study
We conduct a thorough validation of the key components of the SGFormer model under the same experimental conditions. As shown in Table III, initially, we establish a baseline with a model that does not include any specific components, using a Resnet-34 as the encoder and a ViT-based [29] decoder. Then we sequentially introduce core components on the Structured3D and Pano3D datasets to explore the effectiveness of our model.
@@ IV-C1 Effectiveness of curve local embedding (CLE)
From Table III, we can observe that on the Structured3D dataset, there is a 6.70% improvement (Abs.rel) and 6.55% (RMS.lin). On the Pano3D dataset, the performance is further boosted by 1.67% (Abs Rel) and 5.78% (RMS.lin). These results illustrate that by incorporating CLE, we introduce additional spherical surface distance, which aids the network in perceiving the distribution of distortion.
@@ IV-C4 Effectiveness of bipolar re-projection (BRP)
IV-C5 Backbones

### v3

@@ IV-B Comparison Results
Quantitative Analysis: To validate the effectiveness of our approach, we compare our SGFormer model with the current state-of-the-art methods, including strategies for bi-projection fusion and long-range dependencies. Table I shows that our model ranks first in nearly all evaluation metrics, indicating superior overall performance. Specifically, on the one hand, compared with Bifuse [1] based on bi-projection fusion, our model has achieved a 53% improvement (Abs.rel) and 41% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, the performance is significantly improved by 66% (Abs.rel) and 51% (RMS.lin). On the other hand, compared with EGFormer [2] based on long-range dependencies, our model achieves an 11% improvement (Abs.rel) and 12% (RMS.lin) on the Structured3D dataset. For the Pano3D dataset, SGFormer achieves 12% (Abs.rel) and 9% (RMS.lin) gain. In quantitative comparison, we achieve significant performance improvements because we fully leverage the geometric priors of the sphere to relieve the panoramic distortion. The lower quality of depth estimation in SliceNet [13] is mainly due to the detail loss when processing equirectangular images in the vertical direction.
@@
More Comparison Results: We also compare our method with the latest method (CVPR’24) on additional datasets, despite the limitations of the Stanford2D3D [62] and Matterport3D [64] datasets. For instance, the Pano3D dataset [17] incorporates a rendered version of Matterport3D, making the use of both Pano3D and Matterport3D for evaluation redundant. Furthermore, in the Stanford2D3D dataset, the top and bottom parts of equirectangular images are not rendered properly, which handicaps our method when tested on them. We strictly follow the Elite360D [57]. The quantitative comparison is shown in Table II. We achieve better performance on the Stanford2D3D dataset [62] and Matterport3D dataset [64]. Our dataset split is consistent with Elite360D [57], following the official split [62, 64]. For the Matterport3D dataset, we take 61 rooms for training and the others for testing. For the Stanford2D3D dataset, we take the fifth area (area 5) for testing, and the others are for training.
TABLE III: Ablation studies of individual components. We begin with a baseline model, adding each component sequentially. We conduct a series of ablation studies on the S3D (Structured3D) and Pano3D datasets. Results in bold indicate the best results. (CLE: curve local embedding, GCPE: global conditional position embedding, CR: circular rotation, BRP: bipolar re-projection)
@@ IV-C Ablation study
We thoroughly validate the key components of the SGFormer model under the same experimental conditions. As shown in Table III, initially, we establish a baseline with a model that does not include any specific components, using a Resnet-34 as the encoder and a ViT-based [36] decoder. Then, we sequentially introduce core components on the Structured3D and Pano3D datasets to explore the effectiveness of our model.
@@ IV-C1 Effectiveness of curve local embedding (CLE)
From Table III, we can observe that on the Structured3D dataset, there is a 6.70% improvement (Abs.rel) and 6.55% (RMS.lin). On the Pano3D dataset, the performance is further boosted by 1.67% (Abs Rel) and 5.78% (RMS.lin). These results illustrate that by incorporating CLE, we introduce additional spherical surface distance, which aids the network in perceiving distortion distribution.
@@ IV-C4 Effectiveness of bipolar re-projection (BRP)
IV-C5 GCPE
From Fig. 9, the depth details can be observed clearly, indicating that the query-based GCPE strategy can contribute to the spatial structure and depth details.
IV-C6 Backbones